The world price of oil depends on the world supply of oil. The supply schedule is as follows:

World Supply (barrels per day) 4 million 6 million 8 million
World price per barrel $25 $15 $10

The country Iran had a marginal cost of $2 per barrel to extract oil. Iraq has a marginal cost of $4 per barrel. Each country can either produce a maximum of 4 million barrels per day, or a minimum of 2 million barrels per day.

a.) If Iran and Iraq decided to collude, compute the price at which they would get the most profits
Answer #1

a) If both the countries decide to collude and then they decide to produce 4 million barrels of oil , both Iran and Iraq will produce 2 million barrels of oil each.

At the production of 4 million barrels of oil, the world price would be $25 per barrel of oil.

So as we know the marginal cost for Iran is $2 per barrel and marginal cost for Iraq is $4 per barrel.

The profits for Iran = ($25-$2)*2 = 23*2 = $46 million.

The profits for Iraq = ($25-$4)*2 = 21*2 = $42 million

So, total profits = $46+$42 = $88 million

If they collide and decide to produce 6 million barrels of oil, thenOne country would produce 2 million barrels of oil and one country will produce 4 million barrels of oil.

One country would produce 2 million barrels of oil and one country will produce 4 million barrels of oil.

If Iran produces 2 million barrels of oil and Iraq produces 4 million barrels of oil the total profit will be =

(15-2)*2 + (15-4)*4 = 26+44 = $70 million

Now if Iran produces 4 million barrels of oil Anurag produces 2 million barrels of oil a total profit will be=

(15-2)*4 + (15-4)*2 = 52+22 = $74 million

If they collude and decide to produce 8 million barrels of oil, then each country will produce 4 million barrels of oil each, the total profits will be =

(10-2)*4 + (10-4)*4 = 32+24 = $56 million

We can see that if both the firms collude and produce 4 million barrels of oil in total and each country produces 2 million barrels of each, the price of oil will be $25 per barrel of oil and this will give them the maximum profit of $88 million.